Suppose we are given a Boolean function f(x, y, z)=(x+y)(x+y)(x+2). Find its DN form

1
Expert's answer
2021-12-13T13:45:42-0500

A logical formula is considered to be in DNF if it is a disjunction of one or more conjunctions of one or more literals.


"f(x, y, z)=(x+y)(x+y)(x+2)=(x+xy+y)(x+2)="


"=x+2x+xy+xy+xy+2y=x+xy+y=x\lor (x\land y)\lor y"